# Efficient Biological Data Acquisition through Inference Set Design

### v2

@@ E.2 Classification task on RxRx3
Results on the RxRx3 dataset demonstrate minor improvements from the active agents (LC and BALD) compared to random selection. The accuracy on inference and test set remains low throughout the experiment regardless of the acquisition method. This is expected considering the extreme difficulty of predicting biological relationships in a low data regime. The result suggests that in the setting with low predictive power, even when the majority of the data is acquired, the ability of inference set design to provide significant budget reductions is limited.
@@ F.1 Empirical validation of weak calibration assumption
Figure 13: Model calibration analysis for a run of the least confidence agent at different active learning steps (columns) for all of our classification experiments (rows). The colored bars represent the accuracy of predictions binned w.r.t their confidence level, and the gray bars show the identity function illustrating what perfect calibration would look like. The models satisfy the condition of being weakly calibrated since the accuracy of the model’s predictions increases monotonically with their confidence. The confidence distribution shifting to the right as $t$ increases indicates that the growing confidence of the model in correctly predicting the labels of the remaining examples in the inference set.
@@ F.2 Empirical validation of stopping criterion
To empirically validate this bound (Equation 5), we run a large number of trials for both the least-confidence (LC) and random agents across the classification datasets used in the experiments section. The results show that for the LC agent, all trials ended at a stopping time $t=\tau$ at which the true system accuracy $\mu_{\text{sys}}^{t}$ was above the threshold $\gamma$, which is in accordance with the fact that the LC agent uses a lower-bound estimate in place of an unbiased estimate for the inference set accuracy, resulting in a looser bound for Equation 5 and an actual failure-probability lower than $\delta$. For the random agent, we observed $1\%$ of the trials end with a system accuracy below the threshold in one of the experiment, $6\%$ in another, and $0\%$ on the three remaining datasets. These results are in accordance with the theory presented in Section 2.2 and Appendix A.2. On QM9, $6$ out of $100$ experiments resulting in slight bound failure is statistically in accordance with the theoretical bound failure probability of $\delta=5\%$. For the other datasets, the bound was looser, which is also a possibility, and showcases that the observed gap between $\delta$ and the true bound-failure probability can be problem-dependent.

### v3

@@ E.2 Classification task on RxRx3
Our results on the RxRx3 dataset demonstrate minor improvements from the active agents (LC and BALD) compared to random selection. The accuracy on the inference and test sets remains low throughout the experiment regardless of the acquisition method. This is unsurprising considering the extreme difficulty of predicting biological relationships in a low data regime. The result suggests that in the setting with low predictive power, even when the majority of the data is acquired, the ability of inference set design to provide significant budget reductions is limited.
@@ F.1 Empirical validation of weak calibration assumption
Figure 12: Model calibration analysis for a run of the least confidence agent at different active learning steps (columns) for all of our classification experiments (rows). The colored bars represent the accuracy of predictions binned w.r.t their confidence level, and the gray bars show the identity function illustrating what perfect calibration would look like. The models satisfy the condition of being weakly calibrated since the accuracy of the model’s predictions increases monotonically with their confidence. The confidence distribution shifting to the right as $t$ increases indicates the growing confidence of the model in correctly predicting the labels of the remaining examples in the inference set.
@@ F.2 Empirical validation of stopping criterion
To empirically validate this bound (Equation 5), we run a large number of trials for both the least-confidence (LC) and random agents across the classification datasets used in the experiments section. The results show that for the LC agent, all trials ended at a stopping time $t=\tau$ at which the true system accuracy $\mu_{\text{sys}}^{t}$ was above the threshold $\gamma$, which is in accordance with the fact that the LC agent uses a lower-bound estimate in place of an unbiased estimate for the inference set accuracy, resulting in a looser bound for Equation 5 and an actual failure-probability lower than $\delta$. For the random agent, we observed $1\%$ of the trials end with a system accuracy below the threshold in one of the experiments, $6\%$ in another, and $0\%$ on the three remaining datasets. These results are in accordance with the theory presented in Section 2.2 and Appendix A.2. On QM9, $6$ out of $100$ experiments resulting in slight bound failure is statistically in accordance with the theoretical bound failure probability of $\delta=5\%$. For the other datasets, the bound was looser, which is also a possibility, and showcases that the observed gap between $\delta$ and the true bound-failure probability can be problem-dependent.